OpenStudy (anonymous):

Im going to try and post this question again and see what happens... hopefully it wont freeze on me... Perform the operation 6 + SR (-18) all divided by 3

OpenStudy (anonymous):

\[6+\sqrt{-18}/3\]

OpenStudy (anonymous):

i believe you would simplify the 6 to 2 and the 3 to a 1. but for the sqrt i don't know if you have to use the imaginary number (i) or not to simplify it.

OpenStudy (anonymous):

you could simplifty it to 3x3x2. since you have two threes but them one the outside of the root like so..\[3\sqrt{2}\] but i don't know what to do with the negative..

OpenStudy (anonymous):

that is where my concern is... I dont know what to do with the negitive

OpenStudy (anonymous):

what math are you taking?

OpenStudy (anonymous):

college math 2- elementary and intermediate algebra

OpenStudy (anonymous):

were you the one asking me about intermediate algebra yesterday? but i might have it hold on.

OpenStudy (anonymous):

idk if it was yesterday but I was on here the other day

OpenStudy (anonymous):

im reading that your supposed to use ( i ) to replace the negative. did you learn something with i?

OpenStudy (anonymous):

i havent really learned anything, I am going to college online and its all pretty much "self taught"

OpenStudy (anonymous):

hmm well my guess would be either \[2+3i \sqrt{2}\] or\[2+3\sqrt{2i}\]

OpenStudy (anonymous):

thank you, from what I have gathered from the rest of my assignments, as long as the work is shown no matter what you get the credit... thank you again for your help!

OpenStudy (anonymous):

of course! functions and rations are DIFFICULT!! but i'm gladd i could help :0]